OpenStudy (anonymous):

Can the set of lengths be the side lengths of a right triangle? 18 m, 24 m, 30 m

OpenStudy (anonymous):

for a right triangle, (longest side)^2 = sum of other two sides squared So \[30^{2} should equal 18^{2} + 24^{2}\] 30^2 = 900 18^2 + 24^2 = 900 hence this is a right triangle